OpenStudy (anonymous):

Which is a factored form of 27x^3+64?

OpenStudy (solomonzelman):

\[HINT:~~~~~~~27x^3=(3x)^3~~~~~and~~~~~64=4^3\]\[Also:~~~~~~a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

Okay so is my answer (3x+4)(9x^2-12x+16)?

OpenStudy (solomonzelman):

Yes!

OpenStudy (anonymous):

Thank you!